OpenStudy (anonymous):

Help please :0 I need to solve for x log base 3 (10) - (1/log base x(3)) = 3 Any help appreciated just a little lost :(

Nnesha (nnesha):

\[\huge\log _{3} 10 - \frac{ 1 }{ \log }_{x} 3 =3\] hmmm is it your question ??

OpenStudy (anonymous):

the 3 is under with log base x

OpenStudy (anonymous):